OpenStudy (anonymous):

solve the equation. cos x(cos x - 1) = 0

OpenStudy (psymon):

Set each factor, independently, equal to 0 and solve for x cosx = 0 cosx = 1 Now its just knowing which values of cosine equal 0 and 1.

OpenStudy (anonymous):

Umm I still don't get it lol

OpenStudy (jdoe0001):

=> cos x(cos x - 1) = 0 # solve for cos(x) , what do you get?

OpenStudy (anonymous):

1 or 0??

OpenStudy (jdoe0001):

well, yes, as @Psymon said, now look at your Unit Circle, where's cosine 0 and 1? those are very well known angles

OpenStudy (anonymous):

x=0, 90, 180, 270,360,450, 540,630

OpenStudy (anonymous):

0 and 180?

OpenStudy (anonymous):

x=k*90 where k is a integer

OpenStudy (anonymous):

ummm so do I just pick one of the x values to substitute for k?

OpenStudy (jdoe0001):

well, at 180 cosine is -1, not 1

OpenStudy (jdoe0001):

\(\bf cos(x) = 1 \implies cos^{-1}(cos(x)) = cos^{-1}(1)\implies x = 0^o\)

OpenStudy (jdoe0001):

but wait! there's more! -- taken from late night tv shows selling stuff--

OpenStudy (anonymous):

lol oh?

OpenStudy (jdoe0001):

0 degrees is only 1 revolution, that is 2pi if you go around in circles, 4pi, 6pi, 8pi, 64pi .... so the answer is like \(\bf x = n + 2\pi\) where "n" is an integer \(\ge 0\)

OpenStudy (jdoe0001):

well, I guess it should be \(\bf x = 2\pi n\) rather

OpenStudy (jdoe0001):

then you'd have to accomodate the other angle in the notation

OpenStudy (anonymous):

sorry for mistake,

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=cos%28x%29*%28cos%28x%29-1%29%3D0 you can find easy solutions here to all of ur equation